# supplemental

### v1

@@ Abstract
Searching is a task that arises in numerous domains. In nature, examples range from proteins locating their binding sites within the cell [1] to foraging by macro-organisms [2, 3]. In computer science, search algorithms have long been of fundamental interest [4] and have gained cultural importance in determining how the large and unstructured body of data that constitutes the internet is experienced by users [5]. Crossovers between these domains also exist, such as biologically-inspired optimization algorithms which can be applied to solve many and varied science and engineering problems [6]. A large literature has established optimal random search strategies, in the sense of optimizing the efficacy of the search [7, 8, 9, 10, 11].
@@
In this work, we study a feature that is common to all such search processes, namely that increasing efficacy typically comes with a cost, for example, the amount of time or energy that must be invested or the complexity of the algorithm. A natural question is whether one can identify a point of diminishing return, i.e., a point beyond which investing more effort into the search is not compensated by sufficiently increased success. We answer this question by introducing a quantity called profligacy that expresses the cost-efficacy trade-off in a manner similar to Helmholtz free energy, wherein the search cost plays the role of energy, the success probability provides an analog to entropy, and a temperature-like quantity $\lambda$ has the units of cost and characterizes how much effort a searcher is prepared to invest. We will show below that in the context of even fairly simple diffusive searches [12, 13, 14], the optimal strategy that arises from minimizing profligacy exhibits a rich phase diagram (presented in Fig. 2 below). By this we mean that one finds both continuous and discontinuous transitions between regimes in which the optimal search strategy switches. Moreover, the phase structure goes beyond what is normally seen at equilibrium, exhibiting re-entrant behavior and overhangs, to be detailed below.
We begin by deriving the profligacy $\xi$. Its origin is most clearly understood by appealing to an ensemble of $K$ independent searchers, all of whom are following a strategy that is controlled by a parameter $r$. Associated with each searcher is a non-negative cost $C$ that depends on the path that they have followed up to the allotted search time, $t_{f}$. Meanwhile, the overall efficacy of the search can be quantified in terms of the fraction of searchers that have located the target before $t_{f}$. As $K\to\infty$ this fraction converges to $P_{\rm s}(r)$, the success probability for a single searcher (averaged over all possible search histories). Similarly, the total cost (across the ensemble) behaves asymptotically as $K\langle C(r)\rangle$, where $\langle C(r)\rangle$ is the mean cost for a single searcher (again averaged over all possible search histories). To implement the trade-off, we introduce a weighted efficacy $P_{\rm s}(r)\omega(r)$ where the multiplicative factor $\omega(r)$ decreases with the total cost of the search. A natural choice is the exponential function $\omega(r)=\exp(-\langle C(r)\rangle/\lambda)$, since this lies in $[0,1]$ for $\lambda\geq 0$, and we can identify the single parameter $\lambda$, which we call the investment, as a characteristic cost that each searcher is willing to invest. The point of diminishing return can now be defined as the value of $r$ that maximizes the cost-weighted efficacy, or equivalently, that minimizes the profligacy
@@
It is here that we recognize the structure of a Helmholtz free energy that was described above, and within which the variational parameter $r$ relates to the search strategy.
@@
We devote the rest of this work to determining how the search strategy that minimizes the profligacy changes as we vary the investment $\lambda$, within the framework of diffusive searches under stochastic resetting. In this context the searcher is modeled as a diffusive particle starting from the origin, which is reset instantaneously to the origin with rate $r$ [12]. The search is successful if the searcher reaches a target located at distance $m$ from the origin. Early studies of such processes [12, 13, 15, 14] demonstrated how resetting can allow the target to be found more quickly than through diffusion alone, replacing an infinite mean time to locate the target with some finite value. Moreover, there exists an optimal resetting rate $r^{*}$, which minimizes the mean time to find a target [12] and the value of $r^{*}$ undergoes phase transitions as various control parameters are varied [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29]. We emphasize that the optimization problem here is different, in that we seek to minimize the profligacy of a search constrained to end at a predetermined time $t_{f}$. Recently, the consequences of associating a cost with each reset, accounting for the consumption of time, fuel or some other finite resource, have been investigated [30, 31, 32, 33, 34] and the statistics of the cost as a function of $r$ have been computed [33]. Here we consider a predetermined cost of the search, for example, where one purchases a search of duration $t_{f}$ regardless of whether the target is found within that time. Thus the average cost $\langle C(r)\rangle$ does not depend on the target position (see SM [35] for further details). The resetting rate $r$ furnishes a single key parameter, characterizing the search strategy, with which we may optimize the profligacy. We will focus on the transitions from a non-resetting optimal strategy, $r=0$, to an non-zero optimal value of $r$.
@@
We consider two classes of resetting models which differ in what happens at the end time $t_{f}$, as illustrated in Fig. 1. In the case of Resetting Brownian Motion (RBM) [12], the process is simply halted at $t_{f}$, meaning that searcher can be anywhere in space at time $t_{f}$. By contrast, a Resetting Brownian Bridge (RBB) [36] imposes the additional constraint that a searcher must return to the origin at time $t_{f}$. This models such situations as a rescue helicopter having to return to base to refuel after a prespecified flight time. The RBB ensemble is obtained from that of RBM by retaining only those trajectories that occupy the starting position at the completion time, $t_{f}$. In mechanical terms, this conditioning creates a time-dependent drift on the particle motion along with a resetting rate that diverges as $t_{f}$ is approached, thereby guaranteeing a return to the origin [36].
@@
Each reset $i$ contributes a cost $c_{i}\geq 0$ which depends on the distance travelled to the origin in the reset. The number of resets $N$ that occur up to the fixed time $t_{f}$ is a random variable and fluctuates from trajectory to trajectory. The total cost, $C$, of a trajectory is obtained by summing over all $N$ resets that occur along it:
@@
and the cost is zero if there is no reset in the trajectory. Here, $y_{i}=x_{i}/\sqrt{2Dt_{f}}$ is the rescaled (dimensionless) position just before the reset and the function $c(y)$ is the cost per reset. Similarly, it is convenient to use dimensionless, rescaled variables $R=rt_{f}$ and $M=m/\sqrt{2Dt_{f}}$, which eliminate the values of diffusion constant $D$ and search time $t_{f}$ from the discussion, and leave $R$, $\lambda$ and $M$ as the control variables. We will compare a linear, $c(y)=\sqrt{2}y$, and a quadratic, $c(y)=2y^{2}$, cost per reset, since these are the simplest cases for practical applications, but, as we shall see, yield very different phase diagrams. The linear cost can be motivated as the time required to bring the particle back to the origin at a constant velocity [30, 31, 37]. Likewise, the quadratic cost can represent energy consumption, monetary cost or thermodynamic cost for the particle to reset [38, 39, 34].
Our aim now is to determine the optimal resetting strategy—that is, the value of $R=R^{*}$ that minimizes the profligacy (1)—for a given combination of target position $M$ and investment $\lambda$. To achieve this, we must first evaluate the mean cost $\expectationvalue{C}$ and the probability of finding the target $P_{\rm s}$. The resetting systems we consider have the appealing feature that these quantities can be calculated analytically, following recent progress in leveraging renewal properties of the process [33]. The success probability has been calculated for RBM in [12] and for RBB in [36]. The mean costs are derived in detail in the Supplemental Materials [35].
@@
Figure 2: Phase diagrams: heat map of $R^{*}$ which minimizes profligacy $(\xi)$ for 4 different cases: (a) RBM with linear cost (b) RBM with quadratic cost (c) RBB with linear cost and (d) RBB with quadratic cost. The phase boundaries delineate the boundary between regions of zero and non-zero $R^{*}$: a full line indicates a continuous transition and broken line a discontinuous transition. In (a) the horizontal red line indicates the threshold, $M_{\rm T}$, above which there is no transition.
@@
In Fig. 2 we present the phase diagrams in the $\lambda$$M$ plane obtained by minimizing the profligacy with respect to $R$. We distinguish between the two different types of search (RBM and RBB) and the two different cost functions (linear and quadratic). In the unshaded regions, diffusing without resetting is optimal ($R^{*}=0$), whereas in the shaded regions, a nonzero resetting rate yields the least profligate search. Along the solid lines, the optimal resetting rate $R^{*}$ changes continuously across the phase boundary, whilst along the broken lines the optimal resetting rate jumps discontinuously. The nature of the transition is significant: a continuous transition implies that a gradual introduction of resetting yields the optimal search strategy whereas a discontinuous transition implies a sudden switch of strategies to a finite resetting rate.
@@
The easiest phase diagram to understand is that for RBM and a linear cost per reset (Fig. 2(a)) for which the cost (3) increases monotonically with the resetting rate (see SM Fig. S1). When the target is far from the origin (large $M$), the success probability (5) monotonically decreases (see SM Fig. S2). Thus for large $M$, $R^{*}=0$ is always the optimal value. However for $M$ below a threshold value $M_{T}$, the success probability initially increases with $R$ and has a peak at some intermediate value of $R$. This implies that for sufficiently large $\lambda$ the optimal resetting rate $R^{*}$ is non zero. The transition to an optimal strategy that involves resetting can be understood by appealing to a Landau-like theory which implies a classical continuous phase transition into a resetting phase ($R^{*}>0$) as $\lambda$ is increased at fixed $M<M_{T}$.
We now turn to the case of a quadratic cost per reset, for which the mean cost no longer grows without bound as $R\to\infty$ but approaches a plateau (see SM Fig. S1). The effect of this on the RBM phase diagram, Fig. 2(b), is the addition of a discontinuous transition line at intermediate $M$. Although this meets the continuous transition line at two points, it does not end there (as at tricritical points) but extends beyond them, thus creating overhangs (see insets). So as $\lambda$ is increased, for particular choices of $M$, we find an initial continuous transition to a nonzero optimal resetting rate $R^{*}$ very closely followed by a discontinuous jump in $R^{*}$ as shown in Figs. 3(a) and 3(b). This sequence of transitions implies an initial gradual introduction of resetting into the optimal search strategy then a sudden jump to a stronger resetting strategy.
In contrast to RBM, the success probability for a RBB search is a peaked function of $R$ for all target positions $M$ (see SM Fig. S2). The effect of this is that resetting always becomes beneficial for high enough $\lambda$. With a linear cost per reset (Fig. 2(c)), the transition is continuous for high $M$ and discontinuous at low $M$, the two lines meeting at a classical tricritical point (see Fig. 3(c)) [29]. Finally, for the case of RBB with quadratic cost per reset (Fig. 2(d)), the shape of the phase boundary has developed a kink in comparison to the linear cost case. Similar to the case of RBM, the mean cost plateaus as $R\to\infty$, creating an overhang instead of a tricritical point. This overhang effect is evident in Fig. 3(d), where we see both a continuous transition from $R^{*}=0$ to $R^{*}>0$ and a discontinuous jump between two nonzero values of $R^{*}$ at a higher value of $\lambda$. As $M$ is increased, the jump in the discontinuous transition goes to zero and we are left with a single continuous transition. If we consider fixing $\lambda$ and increasing $M$, then in the region of the kink we have re-entrant behaviour into the $R^{*}=0$ phase via a discontinuous transition.

### v2

# Minimizing the Profligacy of Searches with Reset

@@ I Introduction
The task of searching arises in numerous domains. In nature, examples range from proteins locating their binding sites within the cell [1] to foraging by macro-organisms [2, 3]. In computer science, search algorithms have long been of fundamental interest [4] and have gained cultural importance in determining how the large and unstructured body of data that constitutes the internet is experienced by users [5]. Crossovers between these domains also exist, such as biologically-inspired optimization algorithms applied to solve many and varied science and engineering problems [6]. A large literature has established optimal random search strategies, in the sense of optimizing the efficacy of the search [7, 8, 9, 10, 11].
@@
In this work, we study a feature that is common to all such search processes, namely that increasing efficacy typically comes with a cost, for example, the amount of time or energy that must be invested or the complexity of the algorithm. A natural question is whether one can identify a point of diminishing return, i.e., a point beyond which investing more effort into the search is not compensated by sufficiently increased success. We answer this question by introducing a quantity called profligacy that expresses the cost-efficacy trade-off in a manner similar to Helmholtz free energy, wherein the search cost plays the role of energy and the success probability provides an analog to entropy. The arbiter of the competition is a temperature-like quantity $\lambda$, which has units of cost, and characterizes how much extra one is prepared to invest in order to increase the success probability of a search.
We will show below that in the context of even fairly simple diffusive searches, namely Resetting Brownian Motion [12, 13, 14, 15] and Resetting Brownian Bridge [16], the optimal strategy that arises from minimizing the profligacy exhibits a rich phase diagram (presented in Fig. 2 below). By this we mean that one finds both continuous and discontinuous transitions between regimes in which the optimal search strategy switches from a strategy with no resetting to one with a finite resetting rate. Moreover, the phase structure goes beyond what is normally seen at equilibrium, exhibiting re-entrant behavior and overhangs, to be detailed below.
The paper is organized as follows. In Section II we define and motivate the profligacy as a function that captures the cost-efficacy trade-off of a search process. In Section III we define the resetting search processes that we consider. In Section IV we present expressions for the average cost and success probabilities of these searches. In Section V we present phase diagrams obtained by minimizing the progligacy and in Section VI we present a Landau-like theory that explains the features of these phase diagrams. Finally, we conclude in Section VII.
II Profligacy of the search process
In this section we define the profligacy of the search. To be concrete, let us define the set-up we have in mind. Imagine a search company with $K$ independent searchers each of whom is following a search strategy that is controlled by a search parameter $r$, for example it could be Brownian motion with resetting to the origin at rate $r$, see Fig. 1. In addition the search has an allotted time $t_{f}$. Such fixed duration search processes naturally appear in many contexts e.g. when a drone recording data is hired for a fixed time or when a search helicopter with a fixed amount of fuel has a limited flight time to return to its base. Associated with the trajectory of each searcher $j$ is a non-negative cost $C_{j}$. The company also has a budget $C_{b}$ per searcher. For a given realisation of the full search process (trajectories of all $K$ searchers), this budget implies a hard constraint $\sum_{j=1}^{K}C_{j}/K=C_{b}$. We assume that $K$ is very large, such that $\sum_{j=1}^{K}C_{j}/K\to\langle C(r)\rangle$, where angled brackets denote the average of the cost over the trajectories of a ‘single’ searcher. In this limit we have an effective single searcher problem.
A fixed budget $C_{b}$ immediately constrains the value of the search parameter to be $r_{1}$ where $\langle C(r_{1})\rangle=C_{b}$. For simplicity, we assume there is a unique value of $r_{1}$, and that all the searchers start from the origin in a one-dimensional setting. Now imagine that we have the task of finding a fixed target at distance $m$ from the origin, see Fig. 1. The search has to complete the allotted time $t_{f}$ even if the target is located before. For example, an automated drone will continue recording data, to be analysed afterwards, until $t_{f}$. Also, the limited search time $t_{f}$ means there is a finite probability that a searcher may not find the target. We define $P_{s}(r)$ as the success probability, that is, the probability that a searcher locates the target before $t_{f}$.
The goal is to maximize the success probability by choosing an appropriate value of $r$. However, a fixed budget constraint only allows the value $r_{1}$ and $P_{s}(r)$ is typically not maximal at $r_{1}$, but at $r_{2}\neq r_{1}$. This leads to the dilemma of how to maximize $P_{s}(r)$ and at the same time respect budgetary restrictions. This requires one to soften the hard budgetary constraint $\delta(\langle C(r)\rangle-C_{b})$. A natural way to implement this softening is to replace the delta function by an exponential (as is standard in statistical mechanics when one goes from the microcanonical to canonical ensemble by replacing a delta function by an exponential Boltzmann weight). Hence, we consider a weighted efficacy $P_{\rm s}(r)e^{-(\langle C(r)\rangle-C_{b})/\lambda}$ where $\lambda$ is a temperature-like softening parameter in analogy with a Boltzmann weight with $\langle C(r)\rangle-C_{b}$ playing the role of energy. Note that the choice of exponential as the softening function is quite natural. In fact this is the standard procedure in statistical mechanics when one goes from microcanonical to canonical ensemble and replaces a delta function constraint by an exponential weight. However, other choices for the softening function are also possible e.g. a Gaussian centred at $C_{b}$. We expect that the results and conclusions will not change qualitatively, although quantitative details will depend on the choice. In the limit $\lambda\to 0$ the hard constraint is recovered and $r$ is restricted to the value $r_{1}$. In the opposite limit $\lambda\to\infty$, the constraint on budget disappears and the optimal parameter value is $r_{2}$ at which $P_{s}(r)$ is maximized. By varying $\lambda$ one interpolates between the two limits.
We call $\lambda$ the investment, as it has units of cost and characterises how much extra a client is prepared to pay to improve the chance of success. The weighted efficacy can then be expressed as $e^{-(\xi-C_{b})/\lambda}$ where the profligacy $\xi$ is given by
@@ II Profligacy of the search process
Now the objective is to find the value of $r$ that minimizes the profligacy. This is somewhat reminiscent of a Helmholtz free energy where $\langle C(r)\rangle$ plays the role of energy while $\ln P_{\rm s}(r)$ plays the role of entropy. The profligacy (1) is a single quantity that allows one to optimize the trade-off between the competing objectives of maximizing efficacy and minimizing cost.
III Resetting Search Strategies
@@
We devote the rest of this work to determining how the search strategy that minimizes the profligacy changes as we vary the investment $\lambda$, within the framework of diffusive searches under stochastic resetting. In this context the searcher is modeled as a diffusive particle starting from the origin, which is reset instantaneously to the origin with rate $r$ [12], see Fig. 1. The search is successful if the searcher reaches a target located at distance $m$ from the origin. Early studies of such processes [12, 13, 17, 15] demonstrated how resetting allows the target to be found more quickly than through diffusion alone, replacing an infinite mean time to locate the target with some finite value. Moreover, there exists an optimal resetting rate $r^{*}$, which minimizes the mean time to find a target [12] and the value of $r^{*}$ undergoes phase transitions as various control parameters are varied [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31]. Recently, the consequences of associating a cost with each reset, accounting for the consumption of time, fuel or some other finite resource, have been investigated [32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42] and the statistics of the cost as a function of $r$ have been computed [36]. We emphasize that the problem we consider here is different, in that we consider a predetermined time and cost of the search. For example, a client purchases a search period $t_{f}$ regardless of whether the target is found within that time. We adopt the fixed-time time search constraint as it provides a broad framework in which different ensembles of trajectories and different cost functions can be considered. In all these cases the resetting rate $r$ furnishes a single search parameter with which we may optimize the profligacy. We will focus on transitions from a non-resetting optimal strategy, $r=0$, to a non-zero optimal value of $r$.
@@
We consider two classes of resetting models which differ in what happens at the end time $t_{f}$, see Fig. 1. In the case of Resetting Brownian Motion (RBM) [12], the process is simply halted at $t_{f}$, meaning that searcher can be anywhere in space at time $t_{f}$. By contrast, a Resetting Brownian Bridge (RBB) [16] imposes the additional constraint that a searcher must return to the origin at time $t_{f}$. This models such situations as a rescue helicopter having to return to base to refuel after a prespecified flight time. The RBB ensemble is obtained from that of RBM by retaining only those trajectories that occupy the starting position at the completion time, $t_{f}$. In mechanical terms, this conditioning creates a time-dependent drift on the particle motion along with a resetting rate that diverges as $t_{f}$ is approached, thereby guaranteeing a return to the origin [16].
IV Cost and Success Probability Calculations
@@
Let us define the class of cost functions that we consider for the resetting searches. Each reset $i$ contributes a cost $c_{i}\geq 0$ which depends on the distance travelled to the origin in the reset. The number of resets $N$ that occur up to the fixed time $t_{f}$ is a random variable and fluctuates from trajectory to trajectory. The total cost, $C$, of a trajectory is obtained by summing over all $N$ resets that occur along it:
@@
and the cost is zero if there is no reset in the trajectory. Here, $y_{i}=x_{i}/\sqrt{2Dt_{f}}$ is the rescaled (dimensionless) position just before the reset and the function $c(y)$ is the cost per reset. Similarly, it is convenient to use dimensionless, rescaled variables $R=rt_{f}$ and $M=m/\sqrt{2Dt_{f}}$, which eliminate the values of the diffusion constant $D$ and search time $t_{f}$ from the discussion, and leave $R$, $\lambda$ and $M$ as the control variables. We will compare a linear, $c(y)=\sqrt{2}y$, and a quadratic, $c(y)=2y^{2}$, cost per reset, since these are the simplest cases for practical applications, but, as we shall see, yield very different phase diagrams. The linear cost can be motivated as the time required to bring the particle back to the origin at a constant velocity [32, 33, 43]. Likewise, the quadratic cost can represent energy consumption, monetary cost or thermodynamic cost for the particle to reset [44, 45, 37]. We note that other kinds of costs could be relevant to a search process, for example those dependent on the time of resets, distance traveled between excursions, or the number of previous resets etc. However, for simplicity we leave other choices for future studies.
Our aim now is to determine the optimal resetting strategy—that is, the value of $R=R^{*}$ that minimizes the profligacy (1)—for a given combination of target position $M$ and investment $\lambda$. To achieve this, we must first evaluate the mean cost $\expectationvalue{C}$ and the probability of finding the target $P_{\rm s}$. The resetting systems we consider have the appealing feature that these quantities can be calculated analytically, following recent progress in leveraging renewal properties of the process [36]. The success probability has been calculated for RBM in [12] and for RBB in [16]. The mean costs are derived in detail in the Appendix A and the details of the computation of success probabilities is discussed in Appendix B.
@@
Figure 2: Phase diagrams: heat map of $R^{*}$ which minimizes profligacy $\xi$ (1) in the $M$$\lambda$ plane for 4 different cases: (a) RBM with linear cost (b) RBM with quadratic cost (c) RBB with linear cost and (d) RBB with quadratic cost. The phase boundaries delineate the regions of zero and non-zero $R^{*}$: a full line indicates a continuous transition and broken line a discontinuous transition. In (a) the horizontal red line (obtained from the Landau-like expansion (9)) indicates the threshold, $M_{\rm T}$, above which there is no transition. The red dots in (b) and (c) (also obtained from (9)) indicate a transition of a different nature from the classical tricritical point, as described in the text. The figures are obtained by evaluating $\xi$ for an equally spaced set of values of $R$ on a grid of $M$ and $\lambda$. $R^{*}$ is the value of $R$ that minimizes $\xi$. If the value of $R^{*}$ goes from 00 to the minimum non-zero $R$ value (chosen as 0.01) between $M$$\lambda$ grid points, the transition is deemed continuous; if the change is 5 times the minimum $R$ value, it is deemed discontinuous.
@@
V Minimizing Profligacy
Having analytical expressions for the average cost and the success probability, we can now write down the profligacy using (1) and the task at hand is to minimize the profligacy. In Fig. 2 we present the phase diagrams in the $\lambda$$M$ plane obtained by minimizing the profligacy with respect to $R$ for the two different types of search (RBM and RBB) and the two different cost functions (linear and quadratic). In the unshaded regions, diffusing without resetting is optimal ($R^{*}=0$), whereas in the shaded regions, a nonzero resetting rate yields the least profligate search. Along the solid lines, the optimal resetting rate $R^{*}$ changes continuously across the phase boundary, whilst along the broken lines the optimal resetting rate jumps discontinuously. The nature of the transition is significant: a continuous transition implies that a gradual introduction of resetting yields the optimal search strategy whereas a discontinuous transition implies a sudden switch of strategies to a finite resetting rate.
@@ V Minimizing Profligacy
The easiest phase diagram to understand is that for RBM and a linear cost per reset (Fig. 2(a)) for which the cost (3) increases monotonically with the resetting rate (see SM Fig. 7). When the target is far from the origin (large $M$), the success probability (5) monotonically decreases (see SM Fig.8). Thus for large $M$, $R^{*}=0$ is always the optimal value. However for $M$ below a threshold value $M_{T}$, the success probability initially increases with $R$ and has a peak at some intermediate value of $R$. This implies that for sufficiently large $\lambda$ the optimal resetting rate $R^{*}$ is non zero. The transition to an optimal strategy involving resetting can be understood by appealing to a Landau-like theory, to be discussed in Section VI, which implies a classical continuous phase transition into a resetting phase ($R^{*}>0$) as $\lambda$ is increased at fixed $M<M_{T}$.
We now turn to the case of a quadratic cost per reset, for which the mean cost no longer grows without bound as $R\to\infty$ but approaches a plateau (see Appendix A Fig. 7). The effect of this on the RBM phase diagram, Fig. 2(b), is the addition of a discontinuous transition line at intermediate $M$. Although this meets the continuous transition line at two points, it does not end there (as at tricritical points) but extends beyond them, thus creating overhangs (see insets). So as $\lambda$ is increased, for particular choices of $M$, we find an initial continuous transition to a nonzero optimal resetting rate $R^{*}$ very closely followed by a discontinuous jump in $R^{*}$ as shown in Figs. 3(a) and 3(b). This sequence of transitions implies an initial gradual introduction of resetting into the optimal search strategy then a sudden jump to a stronger resetting strategy.
In contrast to RBM, the success probability for a RBB search is a peaked function of $R$ for all target positions $M$ (see SM Fig.8). The effect of this is that resetting always becomes beneficial for high enough $\lambda$. With a linear cost per reset (Fig. 2(c)), the transition is continuous for high $M$ and discontinuous at low $M$, the two lines meeting at a classical tricritical point (see Fig. 3(c)) [31]. Finally, for the case of RBB with quadratic cost per reset (Fig. 2(d)), the shape of the phase boundary has developed a kink in comparison to the linear cost case. Similar to the case of RBM, the mean cost plateaus as $R\to\infty$, creating an overhang instead of a tricritical point. This overhang effect is evident in Fig. 3(d), where we see both a continuous transition from $R^{*}=0$ to $R^{*}>0$ and a discontinuous jump between two nonzero values of $R^{*}$ at a higher value of $\lambda$. As $M$ is increased, the jump in the discontinuous transition goes to zero and we are left with a single continuous transition. If we fix $\lambda$ and increase $M$, then near to the kink we have re-entrant behaviour into the $R^{*}=0$ phase via a discontinuous transition.
VI Landau-like Expansion